OpenStudy (idealist10):

Find the curvature of r(t)=2ti+2tj+k.

OpenStudy (idealist10):

@wio

OpenStudy (anonymous):

The curvature is calculated for every point with: \[k=\left| \frac{f'(t)\times f''(t)}{(f'(t))^3}\right|\]

OpenStudy (anonymous):

where f(t) is your r(t)

OpenStudy (idealist10):

I know that r'(t)=2i+2j and r"(t)=0.

OpenStudy (anonymous):

correct

OpenStudy (idealist10):

So how to do r'(t)xr"(t) and (r'(t))^3?

OpenStudy (idealist10):

(r'(t))^3=(2i+2j)^3?

OpenStudy (anonymous):

i think i should have written the equation more precisely, that should be: \[k=\frac{\left|f'(t)\times f''(t)\right|}{\left|f'(t)\right|^3}\]

OpenStudy (anonymous):

got it now?

OpenStudy (anonymous):

wait im confused

OpenStudy (idealist10):

No. How to do abs(r'(t)xr"(t)) and abs((r'(t))^3)?

OpenStudy (idealist10):

@iambatman @wio

OpenStudy (anonymous):

if you have a vactor r(t) = ai + bj + ck, then \(\left| r(t) \right| = \sqrt{a^2 + b^2 + c^2}\)

OpenStudy (idealist10):

Okay~, I got it. Thanks for the help.

OpenStudy (loser66):

for r'(t) x r"(t) , use determinant rule to find it out, then, take norm of the vector product

OpenStudy (loser66):

the same thing with the denominator, that is you take the norm of r'(t), it is |r'(t)| then ^3 it

OpenStudy (idealist10):

What's the determinant rule?

OpenStudy (loser66):

the x in r'(t) x r"(t) is cross product, right? how to do it? using the way you calculate determinant, right?

OpenStudy (loser66):

ok, r(t) = <2t, 2t, k> r'(t) = <2, 2, 0> r"(t) =<0,0,0> hence your numerator is <2,2,0> x <0,0,0> =<0,0,0> no need to find denominator, the curvature =0

OpenStudy (idealist10):

Thank you, @Loser66.